Calculate the pH of a buffer mixture containing 10 ml of 0.1 M CH3COOH and 1 ml of 0.1
M, CH3COONa.
1
Expert's answer
2013-01-30T10:07:00-0500
Equationto calculate pH of a buffer - Henderson Hasselbalch equation:
pH = pKa + log ([A-]/[HA])
Where [A-] = molar concentartion of conjugate base - [CH3COO-] = 0.1 M*0.01 L
[HA] Molar concentration of weak acid - [CH3COOH] = 0.1M *0.001 L
pKa = -log Ka = -log 1.77*10^-5 = 4.75
Substituting:
pH = 4.75 + log ( [0.01]/[0.001])
pH = 4.75 + log 1.0
pH = 4.75 + 1
pH = 5.75